### v1

@@ 3.3.5 Subjects
The target population of interest consists of people involved in software engineering which work with requirements specifications. We used a non-probability sampling approach based on a mix of purposive and convenience sampling baltes2022sampling . In particular, we wanted to select participants who are diverse in terms of the context variables as determined in Table 1, including their experience, education, and software engineering roles. We approached both students participating in RE courses at our respective institutions and practitioners in our collaborators’ network to purposefully diversify the experience and education of our sample. For all other demographic context factors (e.g., SE roles) we had to rely on convenience sampling.
@@
Among the $25$ participants, $5$ reported a high school degree as their level of education, $8$ a Bachelor’s degree, and $12$ a Master’s degree. No participant reported a Ph.D. degree as their highest degree of education. Figure 6 visualizes the distribution of the participant’s experience in the four domains that were relevant to the experimental requirements: aeronautics, telemetry, databases, and open source. For the latter two, results are balanced across the different knowledge levels, while the former two confirm our assumption that all participants had a low-level knowledge of aeronautics and telemetry systems.
@@
The experiment tool—the Google Draw plugin within the Google Document—was unknown to most ($18$ never used it, $6$ rarely, and $1$ from time to time). A total of $16$ participants ($64$%) report having received a form of training in the modeling activity. The modeling experience (never: $1$, rarely: $11$, from time to time: $10$, often: $3$) resembles a normal distribution.
@@ 3.3.7 Data Collection Procedure
Each session started with the first author explaining the general procedure of the experiment and obtaining consent to evaluate and disclose the anonymized data. No participant refused this consent and all data points could be included in the data evaluation procedure. Then, participants were instructed to read the prepared document in order and complete the contained tasks. The document contained all descriptions about the task such that all participants received the same instructions. The first author oversaw all sessions to address technical difficulties, and recorded the minutes each participant spent per period. Ten minutes were estimated per period, but participants were free to allocate their time. In case participants took longer than the scheduled one hour, they completed the task in as much time as they required. Once the task was complete, participants also filled in the questionnaire to provide demographic information on context variables.
@@ 3.3.8 Data Preparation
To evaluate the collected data, we created a code book that characterizes issues in domain models. We developed detection rules for each dependent variable of the resulting product (i.e., missing entity, superfluous entity, missing association, and wrong association). Then, the first author manually evaluated the resulting domain models of each participant using this codebook and recorded all detected issues.
@@
To assess the reliability of the rating, the fourth author of the paper independently recorded issues of three randomly selected participant responses, yielding an overlap of twelve ratings. Since each domain model can contain an arbitrary number of issues of each type, but our dependent variables only model the number of times that an issue type occurred, we consider each rating of a domain model as a vector of dimensions equal to the number of issue types. We then calculated the agreement between two ratings of the same domain model using the cosine similarity between the vectors. The average cosine similarity is 87.23% and represents substantial agreement.
@@ 3.3.9 Frequentist Data Analysis
We performed a frequentist data analysis of the experimental data as in the original experiment femmer2014impact . Since the factor is categorical and all dependent variables are continuous, our statistical method of choice falls between the parametric paired t-test hsu2014paired or the non-parametric Wilcoxon rank sum test wilcoxon1945individual based on the distribution of the variables, which we evaluated using the Shapiro-Wilk test results shapiro1965analysis .
We reject a null hypothesis if the resulting p-value of a statistical test is lower than the significance level $\alpha$. To account for type I errors when performing multiple hypotheses tests targeting the same independent variable, we apply the Bonferroni correction benjamini1995controlling : we considered $\alpha^{\prime}=\frac{\alpha}{m}$ where $\alpha=0.05$ and $m$ is the number of hypotheses tested for each value of the independent variable. For our five families of hypotheses $\alpha^{\prime}=\frac{0.05}{5}=0.01$.
Additionally, we report the effect size dybaa2006systematic using Cohen’s D for the paired student t-test cohen1969statistical and the matched-pairs rank biserial correlation coefficient for Wilcoxon rank sum test king2018statistical .
@@ 3.3.10 Bayesian Data Analysis
    Treatment: The independent variable, which represents the requirements quality defect present in the requirement.
@@
    Context factors: The independent variables, which represent the properties of the participants.
@@
    Confounding factors: The independent variables, which represent the crossover experiment design (e.g., session, period).
@@
Figure 7 shows the DAG of the running example. The treatment, response variable, and context factors correspond to the study variables as outlined in Table 1. The confounding factors represent the influence of the experimental design. The experimental period blocks the learning effect, i.e., the influence on the response variable caused by repeatedly performing the task. The duration blocks the time effect, i.e., the influence on the response variable caused by the amount of time that a participant took for each instance of the task.
@@
Including an independent variable that has an association with another independent variable may introduce a previously blocked confounding effect by opening a backdoor pearl1995bayesian (introducing spurious associations). Consequently, blindly moving forward with all variables may harm the causal analysis. Instead, the so-called adjustment set of variables needs to be selected mcelreath2020statistical in the identification step. A series of four criteria mcelreath2020statistical allows to make an informed selection of variables to include in the final estimation step. This way, we avoid variable bias like colliders which confound the causal effect between the treatment and the response variable.
In the running example, we assume the following causal relation between independent variables. The more experience a participant has in SE or RE, the more likely it is that they have acquired respective domain knowledge (experience in SE/RE $\rightarrow$ domain knowledge). We need to consider this relationship in the next step to avoid attributing impact to the wrong independent variable. For instance, in the running example, we need to distinguish whether experience in SE/RE has a direct influence on wrong association or whether it just has an influence on domain knowledge, which has an influence on the response variable.
@@
    Group-level intercept ($\alpha_{PID}$, where the results of one participant represent one group): the participant-specific mean of connecting an association wrongly, i.e., the within-subject variance of response variables vegas2015crossover modeled via partial pooling ernst2018bayesian
@@
Finally, we perform a stratified posterior prediction to answer our research questions. To this end, we construct a synthetic data set with four data points, one for each value of the main factor variable (i.e., baseline, PV, AP, PVAP). We fix all other independent variables at representative values—i.e., the mean for continuous and the mode for discrete variables. Then, we sampled $6,000$ predictions for each of the four data points. This isolates the effect of the treatment but maintains the uncertainty of the influence of every independent variable encoded in the standard deviation of every predictor coefficient and, hence, more accurately describes the causal relationship between the treatment and the outcome. We compare the $6,000$ predictions of each of the three treatments (PV, AP, PVAP) with the $6,000$ predictions from the baseline (no defect) and count how often the treatment causes a higher, equal, or lower outcome variable. We scale these values to percentages to summarize the effect of the treatment on the outcome variable. This evaluation avoids a point-wise reduction of the results and comparison to an arbitrary significance level as customary in frequentist analyses gren2021possible . Consequently, we abstain from a binary answer to the hypotheses and rather present the more informative distribution of results.
@@ 4 Results
Section 4.1 shows the results of the frequentist and Section 4.2 the results of the Bayesian data analysis. Section 4.3 compares our results to the original study.
@@ 4.1 Frequentist Data Analysis
Missing Actors
PV 0.10 $(0,\infty)$ 0.39
Missing Entities PV 0.25 $(-1,\infty)$ 0.25 $\ll$ 0.01 $(-1,0)$0.79
@@
Table 2: Results of frequentist analysis including the p-value of the hypothesis test (p), confidence interval (CI), and effect size (ES). Statistically significant results in bold (original study $\alpha=0.05$, this replication $\alpha^{\prime}=0.01$).
@@
[…]
@@
[…]
@@ 4.2.1 Causal Model and Adjustment Set
Figure 9 visualizes the DAG, that graphically models our causal assumptions. It is an extension of Figure 7, the running example, including all five dependent variables. To preserve readability of the DAG, we introduce a distributor node. This node substitutes the connections from the source of every incoming edge to the target of every outgoing edge.
@@
The edges represent the same causal reasoning as presented in the running example in Section 3.3.10. We assume that all independent variables (treatments, context factors, and confounding factors) have an impact on all dependent variables. The impact of the requirements quality defect (the three treatments) is the relationship of interest in our analyses. Additionally, to the already described impact of experience in SE on domain knowledge, we assume the following causal relationships:
@@
    Missing entities impact missing associations: If an entity is missing from the domain model, any association involving that entity will also be missing.
@@ 4.2.2 Posterior Predictions
For example, the first two cells in the second row of Table 3 state that using passive voice causes fewer missing entities in $29.6\%$ and more missing entities in $31.5\%$ of all cases. In the remaining $38.9\%$ of all cases, passive voice causes neither more nor fewer missing entities. Given this balance, the effect of passive voice on missing entities is unclear, and there is not enough evidence to reject $H_{0}^{PV\rightarrow E^{-}}$.
[…]
@@ 4.2.3 Marginal and Conditional Effects
4.3 Comparison with original study results
Firstly, we compare the results of the original study femmer2014impact with the results of our replication. We compare their missing objects with our merged response variable missing entities, as 16/17 entities in our requirements represent objects, and only one represents an actor. Regarding $H_{0}^{PV\rightarrow E^{-}}$, we obtain conflicting results carver2010towards , as we reject the null hypothesis while the original study does not. We obtain differing results for $H_{0}^{PV\rightarrow A^{-}}$, as we cannot reject it as in the original study.
@@ 4.3 Comparison with original study results
[…]

### v3

@@ 3.4.5 Subjects
The target population of interest consists of people involved in software engineering who work with requirements specifications. We used a non-probability sampling approach based on a mix of purposive and convenience sampling baltes2022sampling . In particular, we wanted to select participants who are diverse in terms of the context variables as determined in Table 1, including their experience, education, and software engineering roles. We approached both students participating in RE courses at our respective institutions and practitioners in our collaborators’ network to purposefully diversify the experience and education of our sample. For all other demographic context factors (e.g., SE roles) we had to rely on convenience sampling.
@@
Among the $25$ participants, $5$ reported a high school degree as their level of education, $8$ a Bachelor’s degree, and $12$ a Master’s degree. No participant reported a Ph.D. degree as their highest degree of education. Figure 7 visualizes the distribution of the participant’s experience in the four domains [^2] Domain does not exclusively mean application domain, but rather any coherent ontology related to a specific topic. that contribute semantic knowledge to understanding the requirements: aeronautics, telemetry, databases, and open source. For the latter two, results are balanced across the different knowledge levels, while the former two confirm our assumption that all participants had a low-level knowledge of aeronautics and telemetry systems.
@@
The experiment tool—the Google Draw plugin within the Google Document—was unknown to most ($18$ never used it, $6$ rarely, and $1$ from time to time). A total of $16$ participants ($64$%) report having received a form of training in the modeling activity. The modeling experience (never: $1$, rarely: $11$, from time to time: $10$, often: $3$) resembles a normal distribution. We did not discard data from participants who reported having no modeling experience or formal training in modeling given that our experiment included both comprehensive instructions and a warm-up phase as described in Section 3.4.4.
@@ 3.4.7 Data Collection Procedure
Each session started with the first author explaining the general procedure of the experiment and obtaining consent to evaluate and disclose the anonymized data. No participant refused this consent and all data points could be included in the data evaluation procedure. Then, participants were instructed to read the prepared document in order and complete the contained tasks. The document contained all descriptions of the task such that all participants received the same instructions. The first author oversaw all sessions to address technical difficulties and recorded the minutes each participant spent per period. Ten minutes were estimated per period, but participants were free to allocate their time. In case participants took longer than the scheduled one hour, they completed the task in as much time as they required. Once the task was complete, participants also filled in the questionnaire to provide demographic information on context variables.
@@ 3.4.8 Data Preparation
To evaluate the collected data, we created a code book that characterizes issues in domain models. We developed detection rules for each dependent variable of the resulting product—i.e., missing entity, superfluous entity, missing association, and wrong association—and summarized them in a guideline (available in our replication package frattini2024replication ). Then, the first author manually evaluated the resulting domain models of each participant using this guideline and recorded all detected issues.
@@
To assess the reliability of the rating, the fourth author of the paper independently recorded issues of three randomly selected participant responses, yielding an overlap of twelve ratings. Since each domain model can contain an arbitrary number of issues of each type, but our dependent variables only model the number of times that an issue type occurred, we consider each rating of a domain model as a vector of dimensions equal to the number of issue types. We then calculated the inter-rater agreement of the same domain model using the Spearman rank correlation between the vectors. The average cosine similarity is 77.0% and represents substantial agreement. The two raters discussed the remaining disagreement and concluded that they represented acceptable variance in the interpretation of participants’ responses.
@@ 3.4.9 Frequentist Data Analysis
We performed a frequentist data analysis of the experimental data as in the original experiment femmer2014impact . Since the factor is categorical, all dependent variables are continuous, and our samples are dependent, our statistical method of choice falls between the parametric paired t-test hsu2014paired or the non-parametric Wilcoxon signed-rank test wilcoxon1945individual based on the distribution of the variables, which we evaluated using the Shapiro-Wilk test results shapiro1965analysis .
We reject a null hypothesis if the resulting p-value of a two-tailed statistical test is lower than the significance level $\alpha$. To account for type I errors when performing multiple hypotheses tests targeting the same independent variable, we apply the Bonferroni correction benjamini1995controlling : we considered $\alpha^{\prime}=\frac{\alpha}{m}$ where $\alpha=0.05$ and $m$ is the number of hypotheses tested for each value of the independent variable. For our five families of hypotheses $\alpha^{\prime}=\frac{0.05}{5}=0.01$.
Additionally, we report the effect size dybaa2006systematic using Cohen’s D for the paired student t-test cohen1969statistical and the matched-pairs rank biserial correlation coefficient for Wilcoxon signed-rank test king2018statistical .
@@ 3.4.10 Bayesian Data Analysis
    Treatment: The independent variable that represents the requirements quality defect present in the requirement.
@@
    Context factors: The independent variables that represent the properties of the participants.
@@
    Experimental design factors: The independent variables that represent all factors of the crossover experiment design influencing the response variables vegas2015crossover .
@@
Figure 8 shows the DAG of the running example. The treatment, response variable, and context factors correspond to the study variables as outlined in Table 1. The experimental design variables correspond to the factors listed in Table 2. The experimental period blocks the learning effect, i.e., the influence on the response variable caused by repeatedly performing the task. The duration blocks the time effect, i.e., the influence on the response variable caused by the amount of time that a participant took for each instance of the task. Note that the factor tool experience listed in Table 1 is missing from the DAG since we excluded it as explained in Section 3.4.5. Note that the factor sequence listed in Table 2 is missing from the DAG since it is confounded with subject variability (further explained in Section 5.4.1).
@@
Including an independent variable Z that has an assumed causal effect on both the treatment X (i.e., $Z\rightarrow X$) and the outcome Y (i.e., $Z\rightarrow Y$) opens a non-causal path (i.e., $X\leftarrow Z\rightarrow Y$) from the treatment to the outcome pearl1995bayesian . This so-called backdoor path introduces spurious associations. Consequently, blindly moving forward with all variables may harm the causal analysis. Instead, the so-called adjustment set of variables needs to be selected mcelreath2020statistical in the identification step. A series of four criteria mcelreath2020statistical allows to make an informed selection of variables to include in the final estimation step. This way, we avoid variable bias like colliders which confound the causal effect between the treatment and the response variable.
In the running example, we assume the following causal relation between independent variables. The more experience a participant has in SE or RE, the more likely it is that they have acquired respective domain knowledge (experience in SE/RE $\rightarrow$ domain knowledge). We need to consider this relationship in the next step to avoid attributing impact to the wrong independent variable. For instance, in the running example, we need to distinguish whether experience in SE/RE has a direct influence on wrong association or whether it just influences domain knowledge, which influences the response variable.
@@
    Group-level intercept ($\alpha_{PID}$, where the results of one participant represent one group): the participant-specific mean of connecting an association wrongly, i.e., the within-subject variability of response variables vegas2015crossover modeled via partial pooling ernst2018bayesian
@@
Finally, we perform a stratified posterior prediction to answer our research questions. To this end, we construct a synthetic data set with four data points, one for each value of the main factor variable (i.e., baseline, PV, AP, PVAP). We fix all other independent variables at representative values—i.e., the mean for continuous and the mode for discrete variables. Then, we sampled $6,000$ predictions for each of the four data points. This isolates the effect of the treatment but maintains the uncertainty of the influence of every independent variable encoded in the standard deviation of every predictor coefficient and, hence, more accurately describes the causal relationship between the treatment and the outcome. We compare the $6,000$ predictions of each of the three treatments (PV, AP, PVAP) with the $6,000$ predictions from the baseline (no defect) and count how often the treatment causes a higher, equal, or lower outcome variable. We scale these values to percentages to summarize the effect of the treatment on the outcome variable. This evaluation avoids a point-wise reduction of the results and comparison to an arbitrary significance level as customary in frequentist analyses gren2021possible . Rather than providing a binary answer to the hypotheses, we present the more informative distribution of results. However, for the sake of reporting, we consider the distribution of the duration variable skewed if the two percentages differ from the mean (50%) by 10% each and consider the other distributions skewed if the two percentages differ by 10% from each other.
@@ 4 Results
Section 4.1 shows the results of the frequentist and Section 4.2 the results of the Bayesian data analysis. Section 4.3 compares the part of our results that contributes a conceptual replication to the original study. Section 4.4 compares the results from our FDA to the results from our BDA.
@@ 4.1 Frequentist Data Analysis
Table 3 shows the mean and median values of the response variables similar to how they are reported by Femmer et al. femmer2014impact . Note that the results are not directly comparable to those in the original study as both our experimental objects and treatments varied.
Table 3: Mean and median response variable values (reported as mean/median in each cell)
Defect Duration $D$ Missing Entities $E^{-}$ Superfluous Entities $E^{+}$ Missing Associations $A^{-}$ Wrong Associations $A^{\times}$
none 7.38/6.5 0.23/0 0.5/0 0.38/0 0.08/0
PV 6.88/7 0.81/1 0.42/0 0.81/1 0.62/0
AP 7.12/6 1.23/1 0.96/0.5 1.12/1 0.54/0.5
PVAP 7.96/7 1.27/1 0.46/0 1.38/1 0.38/0
@@
Table 4: Results of frequentist analysis including the p-value of the hypothesis test (p), confidence interval (CI), and effect size (ES). Statistically significant results in bold (original study $\alpha=0.05$, this experiment $\alpha^{\prime}=0.01$).
@@
Missing Actors PV 0.10 $(0,\infty)$ 0.39
Missing Objects PV 0.25 $(-1,\infty)$ 0.25
Missing Entities PV $\ll$ 0.01 $(-1,0)$0.79
@@
    $H_{0}^{\text{{PVAP}}\rightarrow E^{-}}$: the co-occurrence of passive voice and ambiguous pronouns impacts the number of missing entities
@@
    $H_{0}^{\text{{PVAP}}\rightarrow A^{-}}$: the co-occurrence of passive voice and ambiguous pronouns impacts the number of missing associations
@@ 4.2.1 Causal Model and Adjustment Set
Figure 10 visualizes the DAG that graphically models our causal assumptions. It is an extension of Figure 8, the running example, including all five dependent variables. To preserve the readability of the DAG, we introduce a distributor node. This node substitutes the connections from the source of every incoming edge to the target of every outgoing edge.
@@
The edges represent the same causal reasoning as presented in the running example in Section 3.4.10. We assume that all independent variables (treatments, context factors, and experimental design factors) have an impact on all dependent variables. The impact of the requirements quality defect (the three treatments) is the relationship of interest in our analyses. In addition to the already described impact of experience in SE on domain knowledge, we assume the following causal relationships:
@@
    Missing entities impact missing associations: If an entity is missing from the domain model, any association involving that entity will also be missing (as already supported by our re-analysis frattini2024second ).
Equally notable are the non-existing associations between nodes, especially between context factors. In our DAG, we only assume an impact of experience in SE/RE on domain knowledge (as explained in Section 3.4.10). We do not assume, for example, a causal relation between education and experience in SE/RE as higher levels of education do not entail more industrial experience or vice versa. Similarly, we assume education to be independent of domain knowledge as most educational programs known to us are domain-independent. The resulting set of associations visualized in Figure 10 corresponds to the authors’ shared beliefs that warrant assuming causal relationships between two variables. While we do not expect every reader to share these beliefs, we hope that the explicit and transparent documentation of our assumptions invites constructive, iterative improvements by challenging them via empirical investigations.
The DAG from the running example in Figure 8 lists the variable duration as an independent variable, while the final DAG in Figure 10 lists it as a response variable. The variable duration takes on two distinct roles depending on the current analysis. In the case where the analysis targets the effects on the duration, it is the sole response variable. In all other cases, it is an independent variable.
@@ 4.2.2 Posterior Predictions
For example, the first two cells in the second row of Table 5 state that using passive voice causes fewer missing entities in $29.6\%$ and more missing entities in $32.5\%$ of all cases. In the remaining $37.9\%$ of all cases, passive voice causes neither more nor fewer missing entities. Given this balance, the effect of passive voice on missing entities is unclear, and there is not enough evidence to reject $H_{0}^{PV\rightarrow E^{-}}$.
Result of Posterior Predictions The following effects are likely given the skewed distribution of posterior predictions: passive voice, ambiguous pronouns, and their co-occurrence cause an increasing number of missing associations. Ambiguous pronouns cause an increasing number of wrong associations. Ambiguous pronouns cause an increasing number of missing and superfluous.
We use an arbitrary threshold of 10% to report notable results in textual form. Refer to Table 5 for the actual, more fine-grained results.
@@ 4.2.3 Marginal and Conditional Effects
Result of Marginal and Conditional Effects Most context factors do not show a significant impact on either the response variables directly or mediate the effect of quality defects. The few context factors that do show an impact are not significant.
4.3 Comparison of original with our Study Results
For the part of our study that serves as a conceptual replication, we compare the results of the original study femmer2014impact and its re-analysis frattini2024second with our results carver2010towards . Regarding $H_{0}^{PV\rightarrow E^{-}}$, we obtain conflicting results from the FDA as we reject the null hypothesis while the original study does not, but consistent results from the BDA, as the distribution of the posterior prediction of missing entities is balanced. We obtain conflicting results for $H_{0}^{PV\rightarrow A^{-}}$ from the FDA as we cannot reject it as in the original study. Our BDA suggests that passive voice has a slight impact on the number of missing associations (25.5% less and 45.0% more likely to miss an association). The result of the BDA does not suppose an effect as strong as the original study, but it does agree with the re-analysis of the original study frattini2024second on a slight impact. Overall, the results of our BDA agree with the reanalyzed results of the original study. The variation of study elements (e.g., experimental subjects and objects) gomez2010replications increases the replicability space within the generalizability space nosek2020replication and identifies those elements as non-influential juristo2011role to the original claim.
Comparison of Studies The results of the frequentist analyses of the original and our study differ. However, the more thorough Bayesian data analysis agrees with the properly re-analyzed original results. Due to the variation of several elements of our study from the original study, the conceptual replication extends the external validity of the original claim that passive voice has only a slight impact on the domain modeling activity.
4.4 Comparison of FDA with BDA Results
@@
The frequentist analysis rejects $H_{0}^{RQD\in\{PV,AP,\text{{PVAP}}\}\rightarrow E^{-}}$, while the Bayesian analysis remains more cautious. The posterior predictions in Table 5 show a tendency towards the treatment having an impact, but with large uncertainty. Additionally, marginal plots of the Bayesian analysis reveal that the primary role, experience with domain modeling, and education impact the dependent variable. Both analyses agree that $H_{0}^{RQD\in\{PV,AP,\text{{PVAP}}\}\rightarrow E^{+}}$ cannot be rejected, though the Bayesian analysis attributes a tendency of causing superfluous entities to ambiguous pronouns.
The frequentist analysis suggests to reject $H_{0}^{RQD\in\{AP,\text{{PVAP}}\}\rightarrow A^{-}}$, i.e., ambiguous pronouns and their coexistence with passive voice influence the number of missing associations. The Bayesian analysis again shows a tendency towards an impact but retains its uncertainty about the effect. Marginal plots instead emphasize the influence of missing entities on the response variable.
The analyses agree on the impact of ambiguous pronouns on the number of wrong associations and suggest to reject $H_{0}^{RQD\in\{AP,\text{{PVAP}}\}\rightarrow A^{\times}}$. Both the large effect size and the skewed distribution of posterior predictions support the existence of a causal effect of ambiguous pronouns on wrong associations in a domain model.
Comparison of Analysis Methods The results of our frequentist analysis differ from our Bayesian analysis: the Bayesian data analysis remains more cautious about several effects suggested by the frequentist analysis. The extended casual model attributes part of the effect on the response variable on other independent variables than the treatment.